OpenStudy (anonymous):

Can someone help me find the value of r in this equation: 2 x pi x r(.5) + 4pi r^2= pi

OpenStudy (anonymous):

Is this what you equation looks like? \(2\pi r(.5) + 4\pi r^2 = \pi\)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

OK. In the first term, the constants can be combined.

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

it would be πr+ 4πr^2 right?

OpenStudy (anonymous):

sorry i'm new to this so i'm not sure how to add the pi sign and such

OpenStudy (anonymous):

You can use the equation editor or learn \(\LaTeX\). The minimal amount of \(\LaTeX\) needed here is easy to learn. What is \(2 \times 0.5\)?

OpenStudy (anonymous):

Nevermind, I'm an idiot and you are correct.

OpenStudy (anonymous):

So you have \(\pi + 4\pi r^2=\pi\) correct?

OpenStudy (anonymous):

pi r

OpenStudy (anonymous):

http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A) For help learning \(\LaTeX\) if you so desire.

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

If you divide the entire equation by \(\pi\), what do you get?

OpenStudy (anonymous):

r+4r^2=1

OpenStudy (anonymous):

At this point, I would move everything to the left side and factor (or use the quadratic formula).

OpenStudy (anonymous):

Unless I've made a stupid mistake, this answer is going to have a radical in it.

OpenStudy (anonymous):

uh oh, this is part of a larger question and r means radius... but the work makes sense so thanks for your help!

OpenStudy (anonymous):

OK. Feel free to ask another question if you can't work it out. :-)

OpenStudy (anonymous):

do you mind if i explain the bigger question? maybe i set this up wrong

OpenStudy (anonymous):

Sounds like a good idea.

OpenStudy (anonymous):

overview: we're working with a vitamin c tablet thats the shape of a right circular cyinder w a hemisphere attached at each end

OpenStudy (anonymous):

one brand of tablets has a total length of 2 cm and a diameter of .5 cm. i found the surface area for it, which i think is \[\pi\]

OpenStudy (anonymous):

now i have a tablet in which the cylinder has an altitude of .5 cm. find the diameter of the hemispheres of the new tablet if the surface area is the same as the first one

OpenStudy (anonymous):

By total length, do you mean height? And what do you mean by altitude? If you could draw a picture using the draw button below, it may make more sense to me.

OpenStudy (anonymous):

i found all that confusing too. total length of the cylinder and hemispheres. altitude means height of the cylinder i think